Two equal billiard balls are in contact on a billiard table. A third equal ball strikes them symmetrically and remains at rest after impact. The coefficient of restitution is

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Fawz,

let third ball approach the two stationary balls with speed v, let coefficient of restitution be e

from symmetry we can find out that the component of momentum normal to approaching ball is conserved

now consider parallel component: mv=2mucos30

mv=mu√3, coefficient of restitution=u/v=1/√3 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Fawz !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR